OpenStudy (anonymous):

Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

OpenStudy (anonymous):

@Sachintha

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@happyrosy

OpenStudy (anonymous):

do you know substitution method?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you already have a= b-2 Substitute that for a

OpenStudy (anonymous):

a -3b = 4 --> b - 2 -3b = 4

OpenStudy (anonymous):

-2b - 2 = 4 -2b = 6 b = -3

OpenStudy (anonymous):

okay so is it B

OpenStudy (anonymous):

Substitute b in any of your 2 original equations

OpenStudy (anonymous):

No it isn't b. Remember it is in the form of (a, b)---> (a, -3)

OpenStudy (anonymous):

Answer is D i believe

OpenStudy (anonymous):

THANKS! @claringleek

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

a − 3b = 4 a = b − 2 We are going to have to substitute -3 for b. Let's choose the second equation because it is shorter. a = -3 - 2, which means a = -5 So, ( -5, -3) is your answer which is D